OpenStudy (anonymous):

State the missing steps and reasons to this solution of 3(x+4)=18 a) 3(x+4)=18 b) Distributive property c) 3x+12-12=18-12 d) 3x+0=18-12 e) 3x=18-12 f) Simplify g) 3x/3= 6/3 h) 1x=6/3 i) x=6/3 j) x=2

OpenStudy (anonymous):

I know how to do b, c, and g. I need help finding the missing parts. This is not for a test; this is just some warm up problem from school.

OpenStudy (anonymous):

b) 3x+12=18 f)3x=6 Is that what you're looking for? Please ask if you need an explanation :)

OpenStudy (anonymous):

I need help with finding the missing steps and reasons for a, d, e, f, h, i, and j.

Miracrown (miracrown):

\(\color{blue}{\text{Originally Posted by}}\) @padderz b) 3x+12=18 f)3x=6 Is that what you're looking for? Please ask if you need an explanation :) \(\color{blue}{\text{End of Quote}}\) That's right! Good job :)

OpenStudy (anonymous):

I need help with finding the missing steps and reasons for a, d, e, f, h, i, and j.

Miracrown (miracrown):

So you need the reasons for c, d, e, g,h, i and j?Let's start with c a is nothing, a is the starting point

OpenStudy (anonymous):

A is just to write the equation.

Miracrown (miracrown):

Yep! So what are your thought for c?

OpenStudy (anonymous):

Subtract 12 from both sides.

OpenStudy (anonymous):

How about f?

Miracrown (miracrown):

\(\color{blue}{\text{Originally Posted by}}\) @AppleBVB Subtract 12 from both sides. \(\color{blue}{\text{End of Quote}}\) Right!

OpenStudy (anonymous):

What is d?

Miracrown (miracrown):

For f, 3x=6 simplify the right side

OpenStudy (anonymous):

What is d?

Miracrown (miracrown):

Simplify the left side 12-12

OpenStudy (anonymous):

Oh, I get it. So, what about h?

Miracrown (miracrown):

Simplify the division 3/3

OpenStudy (anonymous):

How about i?

OpenStudy (anonymous):

For i, is it to divide by 6.

Miracrown (miracrown):

1.a=a indentity property for the multiplication

OpenStudy (anonymous):

What's j?

Miracrown (miracrown):

Simplify the right side 6/3=2

OpenStudy (anonymous):

Did I forget to ask you "D"?

Miracrown (miracrown):

Nope

Miracrown (miracrown):

Simplify the left side 12-12

OpenStudy (anonymous):

Sorry to be annoying, but what is "e"?

Miracrown (miracrown):

You aint annoying, its all good! For e, indentity element for addition 0+a=a for any a

OpenStudy (anonymous):

Thank you, for our help.

OpenStudy (anonymous):

I mean Thank you, for your help.

Miracrown (miracrown):

Haha, you helped yourself too!